Circular No. 9042

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA IN THE LARGE MAGELLANIC CLOUD 2009 No. 2 W. Liller, Viña del Mar, Chile, reports the discovery of an apparent

nova (mag ≈ 12.1) in the Large Magellanic Cloud from two photographs taken with Kodak Technical Pan film on May 4.994 UT, the position given as α = 5^{h}31^{m}28^{s} ± 3^{s}, δ = −67^{o}05′38′′ ± 8′′ (equinox 2000.0). No star brighter than mag ∼ 19 appears at this position on a U.K. Schmidt image of the field, and nothing brighter than mag 14.0 was visible on Liller’s photograph from 2009 Apr. 23.01. L. A. G. Monard, Pretoria, South Africa, reports a confirmation of the presumed nova at unfiltered mag 12.5 on CCD images taken under less-than-ideal conditions on May 7.716, the position end figures measured as 26^{s}.42, 39′′.4, adding that no star is visible at this position on the Digitized Sky Survey (limiting red mag 20), though there is a star of red mag ≈ 18.5 located ∼ 1′′ southwest of the new object.

2009 HC_{82}

Another minor planet in a retrograde orbit but showing no apparent cometary features has been discovered (cf. MPEC 2009-J04) by the Catalina Sky Survey. The following orbital elements are from MPO 157545:

Epoch = 2009 June 18.0 TT

T = 2008 Nov. 3.8564 TT  ω = 296.^{o}5814
e = 0.810058  Ω = 294.6157 }2000.0
q = 0.463311 AU  i = 154.5642

a = 2.439225 AU n^{o} = 0.2587178 P = 3.81 years

COMETS C/2009 A2, C/2009 A3, C/2009 A4, C/2009 A5 (SOHO) Further to IAUC 9039, additional Kreutz sungrazers have been found

on SOHO website images, being tiny and stellar in appearance (peak mag ∼ 7.5, except C/2009 A2 at mag ∼ 7) in C3 images. C/2009 A3 was small and “roughly stellar” in C2 images. All except C/2009 A2 (condensed, and teardrop-shaped) were slightly diffuse and elongated (peak mag ∼ 11, except C/2009 A2 and A3 at mag ∼ 10) in STEREO HI1-A images.

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 A2 Jan. 1.054 18^{h}55.^{m}8 −25^{o}55′ C3  HS 2009-F18
C/2009 A3  1.613 18 58.2 −25 28 C3/2 HS 2009-F30
C/2009 A4  2.154 19 01.1 −25 14 C3  HS 2009-F30
C/2009 A5  2.488 19 05.1 −25 39 C3  HS 2009-F30
2009 May 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green